The probability that a student has a Visa card (event V) is .68. The probability that a student has a MasterCard (event M) is .21. The probability that a student has both cards is .03.

(a) Find the probability that a student has either a Visa card or a MasterCard. (Round your answer to 2 decimal places.)

Probability


(b) In this problem, are V and M independent?

Yes
No

Answer by ikleyn(52800)

.
(a)  According to the general formula of the Probability theory,


         P(V U M) = P(V) + P(M) - P(V & M) = 0.68 + 0.21 - 0.03 = 0.86.     ANSWER



(b)  According to the general definition of the Probability theory, the events  V  and  M  are independent IF AND ONLY IF  


        P(V & M) = P(V) * P(M).


     In our case IT IS NOT SO ==============>  therefore, the events  V  and  M  are NOT INDEPENDENT.


Solved, answered, explained and completed.